OpenStudy (anonymous):

How would I determine if these two lines are parallel or not? 6x+7y=-28 and y=6/7x-8?

OpenStudy (firekat97):

Find both their gradients

OpenStudy (firekat97):

so put the first one so that y = ...... so 6x + 7y = -28 y = (-28 - 6x)/7 y = -4 - (6/7)x so gradient of this line is -6/7 whereas with the other line, you can read the gradient straight off as 6/7 so since both gradients are not equal (one is positive the other is negative), they are not parallel lines.